If I invested $5,000 at 5% compounded monthly. What is the amount after 28 years? What compound interst formula would I use?
Answer by Theo(13342)

annual interest rate = 5%/100% = .05

monthly interest rate = .05/12 = .004166667

number of time periods = 28 years * 12 compounding periods per year = 336

your formula is:

f = p * (1+i)^n where:

f = future value
p = present amount
i = interest rate per time period (months)
n = number of time periods (months)

this formula becomes:

f = 5000 * (1+(.05/12))^(28*12) which equals:

f = 5000 * (1.004166666666)^(336) and your answer is:

f = $20,217.11097